OpenStudy (dmezzullo):

ind 3x - y - 3z if x = -2, y = 1, and z = -2. -13 -1 1 13

OpenStudy (anonymous):

I don't understand your question. Could you clarify what you are asking?

OpenStudy (dmezzullo):

Evaluate the following expression using the values given. Find 3x - y - 3z if x = -2, y = 1, and z = -2. -13 -1 1 13

OpenStudy (anonymous):

just plug in those values for x, y, z and evaluate the expression.

OpenStudy (swissgirl):

Ok so basically plug into your equation Replace your x with -2 Replace your y with 1 Replace your z with -2 3(-2)-1-3(-2)=-1

OpenStudy (swissgirl):

ohh whoops didnt mean to give the answer lol. Just doing too many things at once here :\